OpenStudy (gahm8684):

@JoannaBlackwelder Given the following perfect square trinomial, fill in the missing term. (Do not type the variable in the blank.) 4x^2 + ___x + 49 Numerical Answers Expected!

OpenStudy (gahm8684):

dont give me the answer only need an explanatipon of how to do it

OpenStudy (joannablackwelder):

Here is the formula: (a + b)^2 = a^2 + 2ab + b^2

OpenStudy (joannablackwelder):

Can you find a and b?

OpenStudy (gahm8684):

but whats a

OpenStudy (joannablackwelder):

You are looking for a single term that is squared.

OpenStudy (joannablackwelder):

what squared is 4x^2

OpenStudy (gahm8684):

(4x+b)^2=4x^2+2(4x^2)b+b^2

OpenStudy (gahm8684):

is it like this

OpenStudy (joannablackwelder):

Not quite.

OpenStudy (joannablackwelder):

(2x)^2 is 4x^2

OpenStudy (joannablackwelder):

So 2x is a.

OpenStudy (joannablackwelder):

Does that make sense?

OpenStudy (gahm8684):

nope:c

OpenStudy (joannablackwelder):

We are using the a^2 + 2ab + b^2 part. Ignore the factored part for now.

OpenStudy (gahm8684):

what factor part;c

OpenStudy (joannablackwelder):

So, the first term (4x^2) is the a^2 term.

OpenStudy (joannablackwelder):

(a + b)^2 This part.

OpenStudy (joannablackwelder):

Sorry, I meant the factored part of the formula.

OpenStudy (gahm8684):

oooh ok i get thanks man

OpenStudy (joannablackwelder):

No worries :) What did you get for the blank?

OpenStudy (gahm8684):

14

OpenStudy (joannablackwelder):

Sweet! Good job!